OpenStudy (anonymous):

Write an equation in slope-intercept form for a line with a y-intercept of 8 and a slope of -1/4. (Points : 2) y = 8x - 1/4 y - 8 = -1/4 x y = -1/4 x + 8 -1/4 x + y = 8

hartnn (hartnn):

the slope intercept form is y=mx+c here, m= slope = -1/4 c= y-intercept = 8 just plug in values! what u get ?

OpenStudy (anonymous):

I haven't gotten anything. I didn't even know where to start

OpenStudy (anonymous):

y=mx+c, is that the formula?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

okay, thank you ! I think i can answer the next one like that

hartnn (hartnn):

good, tell me what answer u get if you want to get your answers verified .